Find the accumulated amount A if the principal P is invested at the interest rate of r/year for t yr. (Round your answer to the nearest cent.)
P = $52,000, r = 7.2%, t = 10.5, compounded quarterly

Answer by rfer(16322)

A=P(1+r/4)^4*10.5
A=52000(1+.018)^42
A=52000(2.115469)
A=$110.004.39